25) Tharaldson Corporation makes a product with the following standard costs:

Standard Quantity or Hours Standard Price or Rate Standard Cost Per Unit
Direct materials 6.5 ounces $ 2.00 per ounce $ 13.00
Direct labor 0.2 hours $ 23.00 per hour $ 4.60
Variable overhead 0.2 hours $ 6.00 per hour $ 1.20

The company reported the following results concerning this product in June.

Originally budgeted output 2,700 units
Actual output 2,800 units
Raw materials used in production 19,380 ounces
Purchases of raw materials 21,400 ounces
Actual direct labor-hours 500 hours
Actual cost of raw materials purchases $ 40,660
Actual direct labor cost $ 12,050
Actual variable overhead cost $ 3,100

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The materials quantity variance for June is:

Multiple Choice

  • $2,360 U

  • $2,242 U

  • $2,360 F

  • $2,242 F
Answer #1

A. $2,360 U

SQ = 2,800 × 6.5 = 18,200

Materials quantity variance = (AQ − SQ) SP

= (19,380 − 18,200) $2 per gram

= $2,360 U